OpenStudy (anonymous):

Please help Which ordered pair is the solution to the system of equations? 2x - y = -5 x + 3y = 22

OpenStudy (anonymous):

A. (1, 7) B. (4, 9) C. (7, 19) D. (8, 21)

OpenStudy (cggurumanjunath):

a

OpenStudy (anonymous):

2x - y = -5 x + 3y = 22 I'm going to eliminate y so I can solve for x. Multiply the top equation by 3. 6x - 3y = -15 x + 3y = 22 Add together. 7x + 0 = 7 7x = 7 x = 1 <--- Plug x into one of the original equations. 2x - y = -5 2(1) - y = -5 2 - y = -5 -y = -7 y = 7 The ordered pair that is a solution to the system of equations is: (1, 7) :D

OpenStudy (cggurumanjunath):

:)

OpenStudy (anonymous):

can you help with one more?

OpenStudy (anonymous):

Yes. (: